The standing wave is formed in a string with two f
Answer #1

a )

given length = 8 m

mass m = 20 g = 0.02 kg

tension T = 40 N

here given third harmonic so \lambda / 2 = L / 3

\lambda / 2 = 8 / 3

\lambda / 2 = 2.666

\lambda = 5.33 m

so the nodes are at 0 , 2.666 , 5.3333 , 8 m

the antinodes will occur halfway between each pair of adjacent nodes

so = 5.333 + 8 / 2 = 6.66 m

b )

we have equation f = v / \lambda

but \mu = m / L

\mu = 0.02 / 8

= 0.0025 kg /m

and v = ( T / \mu )1/2

= ( 40 / 0.0025 )1/2

v = 126.49 m/sec

f = v / \lambda

f = 126.49 / 5.333

f = 23.71 Hz